OpenStudy (anonymous):

Factor out the negative of the greatest common factor. -8x-40y

OpenStudy (skullpatrol):

What is -8 * 5=?

OpenStudy (anonymous):

-40

OpenStudy (skullpatrol):

$$-8x-40y = -8x-8*5y$$

OpenStudy (skullpatrol):

Ok?

OpenStudy (anonymous):

Can't you do it by 4?

OpenStudy (skullpatrol):

@mathstudent55 Let her do it her self

OpenStudy (skullpatrol):

@marixoxo76 Yes you could do it by 4, but the question asks for the GREATEST common factor, right?

OpenStudy (anonymous):

right

OpenStudy (skullpatrol):

Now use the distributive property of multiplication $$ab+ac=a(b+c)$$

OpenStudy (anonymous):

is it -8(x-5)?

OpenStudy (skullpatrol):

$$ -8x-40y$$ what happened to y?

OpenStudy (anonymous):

oh no it's -8(x+5) got it

OpenStudy (anonymous):

thank ya

OpenStudy (skullpatrol):

What happened to the variable y?

OpenStudy (anonymous):

i thought about that too it didn't make sense to me because you're supposed to include it but when i did it online they said with the y is wrong kinda strange

OpenStudy (skullpatrol):

If it is in the original expression it must be in the simplified form.